OpenStudy (fibonaccichick666):

Hey can someone help me understand the process of showing that binary structures are isomorphic?

OpenStudy (fibonaccichick666):

So I get the steps, step 1: Define \(\Phi\) step 2: Show \(\Phi\) is 1-1 step 3: show that \(\Phi\) is onto step 4: show that it's homomorphic

OpenStudy (fibonaccichick666):

but my question is how do you show steps 2,3?

OpenStudy (amistre64):

1-1 implies that f(a) = f(b) if a=b

OpenStudy (amistre64):

onto means that for some y in the range, there is a x in the domain

OpenStudy (fibonaccichick666):

I get that, I guess I should say step 3

OpenStudy (fibonaccichick666):

I understand the meaning, but how do I show/prove that?

OpenStudy (amistre64):

what did you get form step 1?

OpenStudy (fibonaccichick666):

in my case I'm doing it where phi is given as 2n for the integers and addition to the integers and addition

OpenStudy (amistre64):

so:\[\phi(x)=2x\ ?]

OpenStudy (fibonaccichick666):

yup

OpenStudy (amistre64):

code flop

OpenStudy (fibonaccichick666):

I could still read it haha

OpenStudy (amistre64):

then for some y=2x, can we find an suitable x for all ys?

OpenStudy (fibonaccichick666):

no, there are no fractions in the integers for the odd numbers

OpenStudy (fibonaccichick666):

Is that all I would have to say?

OpenStudy (amistre64):

\[y=\phi(x)=2x\] \[\frac{y}{2}=x\] \[y=\phi(\frac{y}{2})\] then you need a different phi right?

OpenStudy (amistre64):

or is y/2 defined in your domain?

OpenStudy (fibonaccichick666):

well, it wanted me to test that particular phi

OpenStudy (amistre64):

x= 0,1,2,3,4,5,6,7,8 2x/2 = x is definantly in x

OpenStudy (fibonaccichick666):

So I could say that using phi=2n, the integers and addition to the integers and addition are not isomorphic?

OpenStudy (fibonaccichick666):

right, but the other way around it is not, correct? It fails onto right?

OpenStudy (amistre64):

phi is a function, the produces an out put if we can define phi(ab) and phi(a) and phi(b) then we can determine of phi(ab)=phi(a)phi(b) ... the homomorphic property

OpenStudy (amistre64):

every even number can be mapped back to some integer

OpenStudy (fibonaccichick666):

wait, but we are doing it for addition not multiplication?

OpenStudy (fibonaccichick666):

so it'd be the sum of 2(a+b)=2a+2b for the homomorphic part

OpenStudy (amistre64):

first step, define some function: f(x) = 2x is the function you defined is it 1-1? for some a,b Integer; 2a = 2b imples that a=b by cancelation is it onto? y = f(x) = 2x, x=y/2; for some x=y/2: y = f(y/2), therefore its onto

OpenStudy (amistre64):

is it homomorphic? thats when the operation becomes important

OpenStudy (fibonaccichick666):

ok, so for onto, that is where I am confused

OpenStudy (fibonaccichick666):

y=f(x)=2x I get x=y/2 I get ^ how that is always in the integers, I don't understand

OpenStudy (amistre64):